OpenStudy (anonymous):

what is the units digit of 2013^2013

OpenStudy (anonymous):

look out the pattern for 3*3*3*------

OpenStudy (anonymous):

oh... so do you know the answer?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

how did you get the answer?

OpenStudy (anonymous):

look out the pattern of unit digit in the multiplication of 3*3*3*------

OpenStudy (anonymous):

3,9,7,1,3,9,7,1.....

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

oh.... thank you soo much

OpenStudy (anonymous):

i still have some other questions...

OpenStudy (anonymous):

can you help me with them?

OpenStudy (anonymous):

u can post?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok...post them

OpenStudy (anonymous):

chris and sandy ran a half mile. sandy ran at the race at a steady pace of 1/6 mile per minute,and chris ran at a steady pace of 1/5 mile per minute. how many seconds after chris finish the race did sandy cross the finish line?

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you know the answer?

OpenStudy (anonymous):

Sandy will take 3 minute and and Chris 2.5 minute

OpenStudy (anonymous):

ok... how did you do it?

OpenStudy (anonymous):

so 30 seconds

OpenStudy (anonymous):

both r running half mile so x* 1/6= 1/2 Chris and y*1/5=1/2 Sandy

OpenStudy (anonymous):

thank you... i still have some more

OpenStudy (anonymous):

let me post them...will yiu help me woth those too?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thank you soo much!!!

OpenStudy (anonymous):

yes...post them

OpenStudy (anonymous):

ok... in a regular hexgon what is a ratio of the length of the shortest diagonal to the length of the longest diagonal? express your answer as a commom fraction in simplest radical form.

OpenStudy (anonymous):



OpenStudy (anonymous):

for shortest one use the cosine formula  cos (theta)= (b^2+ c^2- a^2)/ 2 b*c

OpenStudy (anonymous):

so what is the answer?

OpenStudy (anonymous):

u want the answer only?

OpenStudy (anonymous):

what do you mean by that?

OpenStudy (anonymous):

I mean...i should write the answer directly?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

sqrt(3)/2

OpenStudy (anonymous):

ok... i have more...sorry

OpenStudy (anonymous):

post them

OpenStudy (anonymous):

please...no sorry...it's fine

OpenStudy (anonymous):

ok... liz is a student in Ms.Xu's class. Liz says to her classmates, "Of all pairs of students Ms. Xu can choose as class leaders, i am included in one-tenth of those pairs." how many studets are in Ms. Xu's class?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

if y=-x and y \neq 0\] what is the value of x^2013/y^2013?